A train is on its way to deliver its cargo. The train is travelling at 90 km/h along a straight section of track. It has to cross a bridge to reach the destination. A girl is on the bridge, 2/5 of the way from the train side. The girl can just barely escape the oncoming train by running in either direction. How fast, in km/h, can the girl run? (Assume no acceleration of the girl, instantly at her constant speed.)
1
Expert's answer
2018-10-01T13:45:09-0400
Let the bridge length be L:
t=L/v.
We have:
t=(1-2/5)L/u.
Thus,
L/v=(3/5)L/u
u=3/5 v=3/5 90=54 km/h.